Please help me solve this equation
The sum of two consectutive odd number is 32 and their product is 225 what are the two numbers?

Found 3 solutions by ewatrrr, greenestamps, ikleyn:
Answer by ewatrrr(24785)


Hi
 x + (x+2) = 32  x = 15, and (x+2) = 17
And
 x(x+2) = 15x17 = "255"

Wish You the Best in your Studies.


Answer by greenestamps(13198)



No solution!!

Two consecutive odd numbers with a sum of 32 are 15 and 17; the product of 15 and 17 is 255, not 225.

Perhaps you typed the number for the product incorrectly....


Answer by ikleyn(52778)

.
Please help me solve this equation
the sum of two consecutive odd number is 32 and their product is 225. what are the two numbers?
~~~~~~~~~~~~~~~~


Such integer numbers,  as described,  DO  NOT  EXIST.

The problem  HAS  NO  solutions.


Please do not post it again :   This  "problem"  is a  FAKE.